OpenStudy (anonymous):

the condition that the roots of px^2 - px + q = 0 are in the ratio p:q is?

OpenStudy (zepp):

The only way the roots of a quadratic function is 0 is that the vertex is at the origin, (0,0), x-intercept are at 0. Therefore, this must be the simple parabola function, \(f(x)=x^2\)

OpenStudy (zepp):

Now you know what to do :)

OpenStudy (anonymous):

no

OpenStudy (zepp):

Ok, \(f(x)=x^2\) is exactly the same thing as\(f(x)=1x^2+0x+0\), do you agree?

OpenStudy (zepp):

1 is p 0 is q Ratio would be 1:0 :) Got it?

OpenStudy (anonymous):

The answer should be 2p + q =0

OpenStudy (anonymous):

let roots be a and b sum of roots = a+b = 1 and ab = q/p or 1/ab = p/q and you want a/b = p/q so 1/ab = a/b => a=1 or -1 (assuming one of the roots is not 0 i.e. q not equals 0) for a=1, b=0 => not possible so a= -1 and b =2 --> req condition i may be wrong somewhere though,,hmm you may also that p/q should be -1/2

OpenStudy (anonymous):

well I can think that u know that if alpha and beta are roots of the equation then  now i have given u two equations slove